OpenStudy (fibonaccichick666):

Calc III students! Can you give me some difficult integral problems you may have been assigned without answers? (I need to study for the GRE)

OpenStudy (anonymous):

What is in the syllabus?

OpenStudy (fibonaccichick666):

..? There is no syllabus?

OpenStudy (anonymous):

I mean what are the topics in Calc III?

OpenStudy (fibonaccichick666):

Uh essentially, triple integrals, polar, spherical, vector integration, partial derivs. That type of stuff, Idk, really anymore, hence the question. It's been like 4 years since I had it

OpenStudy (zzr0ck3r):

\[\int_{-\infty}^{\infty}e^{-x^2}dx\]

ganeshie8 (ganeshie8):

thats a fun integral :) I think @eliassaab sir has some website to practice calcIII problems

OpenStudy (fibonaccichick666):

...Don't you have to go into the complex plane for that? It's the definition of the zeta function or something right?

OpenStudy (fibonaccichick666):

straight computation yields a divergent integral to infinity, no?

OpenStudy (zzr0ck3r):

no, hint \(\int_a^b f(x) dx = \int_a^bf(y)dy\\x^2+y^2=r^2\)

OpenStudy (fibonaccichick666):

oh god... I have to go to into spherical coordinates?...It really has been a while. I have no idea what your hint means if that is not it

OpenStudy (zzr0ck3r):

no, polar

OpenStudy (fibonaccichick666):

okhm

OpenStudy (anonymous):

\[\int\limits 1+\sin \theta divcos \theta \sqrt{4\cos^2\theta+(1+\sin \theta)^2}d\]

OpenStudy (fibonaccichick666):

\[\int_{-\infty}^{\infty}e^{-x^2}dx\]\[x^2=r^2-y^2\]\[-\infty <y^2<\infty\] \[-\infty <r^2<\infty\]\[\int \int e^{-(r^2-(rsin\theta^2)}rdrd\theta\]

OpenStudy (fibonaccichick666):

the squared sould be on the entire rsintheta term, what do you think?

OpenStudy (fibonaccichick666):

@danish071996 I've never seen divcos, what function is that?

OpenStudy (anonymous):

oh it is division

OpenStudy (fibonaccichick666):

\[\int\limits 1+\frac{\sin \theta}{ cos \theta} \sqrt{4\cos^2\theta+(1+\sin \theta)^2}d\theta\] Is this it?

OpenStudy (fibonaccichick666):

How do we know that e^-x^2=e^-y^2?

OpenStudy (fibonaccichick666):

well, the itegral of those

OpenStudy (anonymous):

no only 1+sin theta in the numerator

OpenStudy (fibonaccichick666):

ah \[\int\limits \frac{1+\sin \theta }{cos \theta \sqrt{4\cos^2\theta+(1+\sin \theta)^2}}d\theta\]

OpenStudy (zzr0ck3r):

\(\int_{-\infty}^{\infty}e^{-x^2}dx=\int_{-\infty}^{\infty}e^{-y^2}dy\\ \int_{-\infty}^{\infty}e^{-x^2}dx*\int_{-\infty}^{\infty}e^{-y^2}dy=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-x^2}e^{-x^2}dx \ dy=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-(x^2+y^2)}dx = \\ \int_0^{2\pi}\int_0^\infty e^{-r}r\ dr \ d\theta = \int_0^{2\pi}\frac{1}{2} d\theta = \pi \) so \( (\int_{-\infty}^\infty e^{-x^2}dx)^2=\pi\) thus \[\int_{-\infty}^\infty e^{-x^2}dx=\sqrt{\pi}\]

OpenStudy (fibonaccichick666):

@zzr0ck3r , can you explain the initial reasoning of f(x)=f(y)?

OpenStudy (zzr0ck3r):

\(\int_2^3x^2dx=\int_2^3y^2dy\)

OpenStudy (fibonaccichick666):

oh, just a change of variable? Bt why do it?

OpenStudy (zzr0ck3r):

well keep reading and it should become clear.

OpenStudy (fibonaccichick666):

unfortunately it cuts off, also wouldn't we have to divide by the integral at the end or take the square root since we are essentially squaring it?

OpenStudy (zzr0ck3r):

\(\int_{-\infty}^{\infty}e^{-x^2}dx=\int_{-\infty}^{\infty}e^{-y^2}dy\\ \implies (\int_{-\infty}^{\infty}e^{-x^2}dx)(\int_{-\infty}^{\infty}e^{-y^2}dy)= (\int_{-\infty}^{\infty}e^{-x^2}dx)^2\)

OpenStudy (fibonaccichick666):

I see that it works quite beautifully into polar though

OpenStudy (fibonaccichick666):

but, yea, that last post is where I do not follow, Is there something in the transition that negates the exponent?

OpenStudy (zzr0ck3r):

\(\int_{-\infty}^{\infty}e^{-x^2}dx=\int_{-\infty}^{\infty}e^{-y^2}dy\\ \int_{-\infty}^{\infty}e^{-x^2}dx*\int_{-\infty}^{\infty}e^{-y^2}dy= \\ \int_{-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-x^2}e^{-x^2}dx \ dy=\\ \int_{-\infty}^{\infty}\int_{-\infty}^{\infty}e^{-(x^2+y^2)}dx = \\ \int_0^{2\pi}\int_0^\infty e^{-r}r\ dr \ d\theta = \\ \int_0^{2\pi}\frac{1}{2} d\theta = \pi \)

OpenStudy (zzr0ck3r):

So what we have shown is \((\int_{-\infty}^\infty e^{-x^2}dx)^2= \int_{-\infty}^{\infty}e^{-x^2}dx*\int_{-\infty}^{\infty}e^{-y^2}dy=\pi\) so \(\int_{-\infty}^\infty e^{-x^2}dx=\sqrt{\pi}\)

OpenStudy (fibonaccichick666):

oh, ohk, I will need to repeat this

OpenStudy (fibonaccichick666):

Where did you find that problem?

OpenStudy (zzr0ck3r):

On here I think.

OpenStudy (fibonaccichick666):

hmm, ok, it's a really good problem to show me I need to review a lot. I would have never thought of that

OpenStudy (fibonaccichick666):

Thank you, I really appreciate it!

OpenStudy (zzr0ck3r):

not many people do, don't worry. It is a very clever trick and I am sure most people that know it have been taught it and did not just come up with it.

OpenStudy (fibonaccichick666):

If you have any others you stumble upon, please post more if you have them. (A few grad students have said the Math GRE is just the hardest calc III integrals essentially, so I want to go through a lot of them in order to avoid a surprise)

ganeshie8 (ganeshie8):

https://www.youtube.com/watch?v=fWOGfzC3IeY

OpenStudy (anonymous):

Use firefox to access my site on http://moltest.missouri.edu/mucgi-bin/calculus.cgi

OpenStudy (fibonaccichick666):

That is amazing, thank you elias!

ganeshie8 (ganeshie8):

it also shows a complete solution for each problem http://gyazo.com/7e02f0160fb6efc51a357fca070dfc21

OpenStudy (fibonaccichick666):

nice!

OpenStudy (anonymous):

YW @FibonacciChick666